You are considering investing in a company that cultivates abalone for sale to local restaurants. Use the following information:

Sales price per abalone = $43.70
Variable costs per abalone = $10.80
Fixed costs per year = $462,000
Depreciation per year = $137,000
Tax rate = 22%

The discount rate for the company is 14 percent, the initial investment in equipment is $959,000, and the project’s economic life is 7 years. Assume the equipment is depreciated on a straight-line basis over the project’s life and has no salvage value.

a.

What is the accounting break-even level for the project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

b.

What is the financial break-even level for the project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

1.
=(462000+137000)/(43.7-10.8)=18206.6869300912

2.
=((959000*14%/(1-1/1.14^7)-137000)/(1-22%)+462000+137000)/(43.7-10.8)=21582.5535744828